(1) The table below shows the calories, fat, and carbohydrates per ounce for three brands of cereal.

Calories Fat Carbohydrates

Cereal Brand A 10 0 11

Cereal Brand B 100 0.2 22.5

Cereal Brand C 130 5.6 19

Total Required 400 6.2 92


(i) Using the information in the table above, write a system of three equations.

(ii) Write a system in matrix form Ax = b.

(iii) Using the Cramer's Rule or Inverse Method, find the amount of each brand of cereal that will give the level of nutrition required. 


1
Expert's answer
2021-11-21T13:32:55-0500

(i)

"\def\arraystretch{1.5}
 \begin{array}{c}
 10x+100y+130z=400 \\
 0x+0.2y+5.6z=6.2 \\
 11x+22.5y+19z=92
\end{array}"

(ii)


"A=\begin{pmatrix}
 10& 100& 130 \\
 0& 0.2& 5.6 \\
11 & 22.5 & 19
\end{pmatrix}, X=\begin{pmatrix}
 x\\
 y \\
z
\end{pmatrix}, B=\begin{pmatrix}
 400\\
 6.2 \\
92
\end{pmatrix},"

"AX=B"

"\begin{pmatrix}
 10& 100& 130 \\
 0& 0.2& 5.6 \\
11 & 22.5 & 19
\end{pmatrix}\begin{pmatrix}
 x\\
 y \\
z
\end{pmatrix}=\begin{pmatrix}
 400\\
 6.2 \\
92
\end{pmatrix}"

(iii)


"\Delta=\det A=\begin{vmatrix}
 10& 100& 130 \\
 0& 0.2& 5.6 \\
11 & 22.5 & 19
\end{vmatrix}"

"=10\begin{vmatrix}
 0.2 & 5.6 \\
 22.5 & 19
\end{vmatrix}-0\begin{vmatrix}
 100 & 130 \\
 22.5 & 19
\end{vmatrix}+11\begin{vmatrix}
 100 &130 \\
 0.2 & 5.6
\end{vmatrix}"

"=-1222-0+5874=4652\not=0"


"\Delta_1=\begin{vmatrix}
 400& 100& 130 \\
 6.2 & 0.2& 5.6 \\
92 & 22.5 & 19
\end{vmatrix}"

"=400\begin{vmatrix}
 0.2 & 5.6 \\
 22.5 & 19
\end{vmatrix}-6.2\begin{vmatrix}
 100 & 130 \\
 22.5 & 19
\end{vmatrix}+92\begin{vmatrix}
 100 &130 \\
 0.2 & 5.6
\end{vmatrix}"

"=-48880+6355+49128=6603"


"\Delta_2=\begin{vmatrix}
 10& 400& 130 \\
 0& 6.2& 5.6 \\
11 & 92 & 19
\end{vmatrix}"

"=10\begin{vmatrix}
 6.2 & 5.6 \\
 92 & 19
\end{vmatrix}-0\begin{vmatrix}
 400 & 130 \\
 92 & 19
\end{vmatrix}+11\begin{vmatrix}
 400 &130 \\
 6.2 & 5.6
\end{vmatrix}"

"=-3974-0+15774=11800"

"\Delta_3=\begin{vmatrix}
 10& 100& 400 \\
 0& 0.2& 6.2 \\
11 & 22.5 & 92
\end{vmatrix}"

"=10\begin{vmatrix}
 0.2 & 6.2 \\
 22.5 & 92
\end{vmatrix}-0\begin{vmatrix}
 100 & 400 \\
 22.5 & 92
\end{vmatrix}+11\begin{vmatrix}
 100 &400 \\
 0.2 & 6.2
\end{vmatrix}"

"=-1211-0+5940=4729"

"x=\dfrac{\Delta_1}{\Delta}=\dfrac{6603}{4652}=1.41939"

"y=\dfrac{\Delta_2}{\Delta}=\dfrac{11800}{4652}=2.53654"

"z=\dfrac{\Delta_3}{\Delta}=\dfrac{4729}{4652}=1.01655"